Circular No. 7703

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2001 (22) 1 The Central Bureau received word of a new apparent satellite of mi-

nor planet (22) Kalliope first from Merline and over half a day later from Margot. Their reports are published below in the order received.

W. J. Merline, Southwest Research Institute (SRI); and F. Menard, Observatoire de Grenoble, report for their collaboration (L. Close, University of Arizona; C. Dumas, Jet Propulsion Laboratory; and C. R. Chapman and D. C. Slater, SRI) the detection of a satellite of (22) on Sept. 2.6 UT from H-band direct imaging with the 3.6-m Canada-France-Hawaii Telescope (+ PUEO adaptive optics system) on Mauna Kea: “The brightness ratios observed are ∼ 4.9 mag in bands J, H, and K′. A 1.3-hr baseline rules out background stars as the source and shows motion consistent with that of a satellite. A search for known background/foreground small bodies with possibly similar motion reveals no candidates brighter than mag 18. We acquired additional observations of the new object over a 3-hr span on Sept. 3. We find the secondary at the following separations and position angles: Sept. 2.5894, 0′′.56, 181^{o}; 2.6301, 0′′.55, 183^{o}; 3.6458, 0′′.28, 317^{o}.”

J.-L. Margot and M. E. Brown, California Institute of Technology, write: “We report the discovery of a satellite to the M-class minor planet (22). The projected separation between primary and secondary was 0′′.51 (1000 km) on H-band images obtained on Aug. 29.6 UT using the adaptiveoptics system at the 10-m W. M. Keck II telescope on Mauna Kea. The brightness ratio was measured to be 23 ± 5, implying a diameter ratio of ∼ 1:5. Confirming observations were obtained at the Keck II telescope on Aug. 31.6.”

SUPERNOVA 2001ed IN NGC 706 M. Migliardi, Cortina d’Ampezzo, Italy, reports his discovery of an

apparent supernova (mag ∼ 14.9) on unfiltered CCD images taken with a 0.5-m telescope on Sept. 2.96 and 3.01 UT in the course of the CROSS program (cf. IAUC 7373). SN 2001ed is located at α = 1^{h}51^{m}51^{s}.08, δ = +6^{o}17′27′′.4 (equinox 2000.0), which is 9′′ east and 20′′ south of the nucleus of NGC 706. A CROSS image taken on 2000 Dec. 28 showed nothing at the position of the new object (limiting mag ∼ 18.5), and it does not appear on Palomar Sky Survey red and blue plates.

2001 September 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green